OpenStudy (anonymous):

a sphere f mass 40 kg is attracted by a sec sphere of mass 15 kg when their centres are 20 cm apart,with a force of 0.1 milligram weight. calculate the value of gravitational constant.

OpenStudy (anonymous):

.....??????????

OpenStudy (jamesj):

What is Newton's Universal Law of Gravitation? That is what you have to use for this problem.

OpenStudy (anonymous):

gravitational force = 0.1 mg-wt. = 0.1 * 10^-6 * 9.8 N = Gm1m2/r^2 m1 = 40 kg m2 =15 kg and r = 0.2 m Put in and find G